# Motivic classes of stacks in finite characteristic and applications to stacks of Higgs bundles

### v1

@@ 3.2. $\lambda$-ring $R_{q}$
takes $F^{m}\operatorname{Mot}_{sch}(\mathbb{F}_{q})[L^{-1}]$ to $F^{2m}R_{q}$, so it is continuous and can be extended to the completion $\overline{\operatorname{Mot}}_{sch}(\mathbb{F}_{q})$. Finally $\overline{\operatorname{Mot}}_{sch}(\mathbb{F}_{q})\cong\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ by Proposition 2.9. ∎
We define a localization of $R_{q}$ as
$R_{q,fin-type}\coloneqq R_{q}[(L^{i}-1)^{-1}|i>0].$
Note that $L^{i}-1$ is invertible in $\overline{R}_{q}$, so the homomorphism $R_{q}\to\overline{R}_{q}$ factors through $R_{q,fin-type}$. Note that $R_{q}\to\overline{R}_{q}$ is injective, so the induced morphism $R_{q,fin-type}\to\overline{R}_{q}$ is injective as well.
@@
There is a commutative diagram:
(3.1) $\operatorname{Mot}(\mathbb{F}_{q})$$\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$$R_{q,fin-type}.$$\overline{\operatorname{Mot}}(\mathbb{F}_{q})$$\overline{R}_{q}$$\scriptstyle{H^{\ast,st}}$$\scriptstyle{H^{\ast}}$
@@
First we want to show that $H^{\ast,st}$ exists. By Proposition 3.7, we have
$H^{\ast}\colon\operatorname{Mot}_{sch}(\mathbb{F}_{q})[\mathbb{L}^{-1}]\to R_{q}.$
Recall that $H^{\ast}(\mathbb{L})=L$ and
$H^{\ast}(\operatorname{BGL}(n))=\frac{1}{(L^{n}-1)(L^{n}-L)\cdots(L^{n}-L^{n-1})}\in R_{q,fin-type}.$
By Theorem 2.6, the image of
@@
is contained in $R_{q,fin-type}$.
Note that $\operatorname{Mot}(\mathbb{F}_{q})\to\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ factors through $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$, so the homomorphism $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})\to R_{q,fin-type}$ making the diagram (3.1) commutative exists. ∎
We also define a map $\operatorname{tr}\colon R_{q}\to\overline{\mathbb{Q}}_{\ell}\text{ given by }[V,\psi]\mapsto\operatorname{tr}\psi.$ We claim that it is a ring homomorphism. It suffices to check for the basis $E_{\lambda}]\in R_{q,i},[E_{\mu}]\in R_{q,j}$:
$\operatorname{tr}\left([E_{\lambda}][E_{\mu}]\right)=\operatorname{tr}\left([E_{\lambda\mu}]\right)=\lambda\mu=\operatorname{tr}\left([E_{\lambda}]\right)\cdot\operatorname{tr}\left([E_{\mu}]\right).$
Note that the homomorphism $\operatorname{tr}$ takes $L$ to $q$ and $L^{i}-1$ to $q^{i}-1\neq 0$, so by the universal property of localizations, it extends to a ring homomorphism $\operatorname{tr}\colon R_{q,fin-type}\to\overline{\mathbb{Q}}_{\ell}.$
Remark 3.10.
For $V=V_{-}\oplus V_{+}$ a super object of $J_{q}$, we can also define $\operatorname{tr}$ as
$V,\psi]\mapsto\operatorname{tr}\psi|_{V_{+}}-\operatorname{tr}\psi|_{V_{-}}.$
Proposition 3.11.
There is a ring homomorphism
$\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})\to\mathbb{Q}$
making the following diagram commutative:
[…]
Note that #\# is a homomorphism of Proposition 3.2.
Proof.
@@
Recall that $\operatorname{tr}(L)=q$, then by Theorem 2.6, the universal property of localization, we see that the lower pentagon of diagram (3.2) commutes.
Finally we want to show that the ring homomorphism $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})\to\mathbb{Q}$ making the diagram (3.2) commutative exists. Since $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$ is a quotient ring of $\operatorname{Mot}(\mathbb{F}_{q})$, it is enough to show that if $A\in\operatorname{Mot}(\mathbb{F}_{q})$ maps to zero in $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$, then #(A)=0\#(A)=0. This follows from the already constructed part of the diagram by diagram chasing. ∎
Recall that in Section 3.2 we defined a pre-$\lambda$ structure on $R_{q}$. The pre-$\lambda$ structure on $R_{q}$ can be extended to $\overline{R}_{q}$ by continuity. We also know that $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ is a pre-$\lambda$-ring by Theorem 2.18.
@@
Let $X$ be a quasi-projective scheme over the finite field $\mathbb{F}_{q}$. Denote by $\lambda_{H}$ the pre-$\lambda$ structure on $\overline{R}_{q}$. Then by Lemma 3.6,
$\lambda_{H}(-H^{\ast}([X]))=\sum_{i}Sym^{i}H^{\ast}([X])(-t)^{i}.$
On the other hand,
$\lambda(-[X])=\sum_{i}[Sym^{i}X](-t)^{i}.$
It suffices to show that
$\lambda_{H}(-H^{\ast}([X]))=\sum_{i}H^{\ast}([Sym^{i}X])(-t)^{i},$
because it implies that $H^{\ast}$ is a homomorphism of pre-$\lambda$-rings by continuity. However by Deligne’s theorem (see [SGA4, Exposé XVII, Théorème 5.5.21]),
$Sym^{i}H^{\ast}([X])=H^{\ast}([Sym^{i}X]).$
Now we know that $\overline{\operatorname{Mot}}(\mathbb{F}_{q})\to\overline{R}_{q}$ is a homomorphism of pre-$\lambda$-rings. ∎
Proof of Theorem 3.3.
@@
Now we want to show that the image of
[…]
is contained in $R_{q,fin-type}$, i.e., we have a commutative diagram
$\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})$$\overline{R}_{q}$$\overline{\operatorname{Mot}}^{\lambda-ring}_{fin-type}(\mathbb{F}_{q})$$R_{q,fin-type}.$$\scriptstyle{H^{\ast,\lambda-ring}}$
By Proposition 3.9, for any $A\in\operatorname{Mot}(\mathbb{F}_{q})$, we have $H^{\ast,st}(A)\in R_{q,fin-type}$.
@@
We want to show that the ring homomorphism $\overline{\operatorname{Mot}}^{\lambda-ring}_{fin-type}(\mathbb{F}_{q})\to\mathbb{Q}$ making the above diagram commutative exists. Note that $\overline{\operatorname{Mot}}^{\lambda-ring}_{fin-type}(\mathbb{F}_{q})$ is a quotient ring of $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$, then it is enough to show that if $A\in\operatorname{Mot}(\mathbb{F}_{q})$ maps to zero in $\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})$, then #(A)=0\#(A)=0. This follows from the already constructed part of the diagram by diagram chasing. ∎
@@ 4. Power structures
Our goal in this section is to define a power structure on $\overline{\operatorname{Mot}}(k)$ and interpret certain powers as configuration spaces, see Proposition 4.4. These are similar to [GZLMH13, Definition 1] and [BM15, Lemma 6] respectively.
@@ 4.1. From pre-$\lambda$-rings to plethystic exponents and plethystic logarithms
In this section, for an arbitrary pre-$\lambda$-ring, as a generalization of [GZLMH04, Remark 2, Page 55] we define the plethystic exponents and plethystic logarithms in the multivariable case.
Let $A$ be a pre-$\lambda$-ring with the pre-$\lambda$ structure $a\mapsto\lambda_{a}(t)$. Let $A[[z_{1},z_{2},\ldots,z_{n}]]^{+}$ denote the ideal of power series with vanishing constant term and we denote by $(1+A[[z_{1},z_{2},\ldots,z_{n}]]^{+})^{\times}$ the multiplicative group of series with constant term equal to $1$, we define the plethystic exponent
$\operatorname{Exp}\colon A[[z_{1},z_{2},\ldots,z_{n}]]^{+}\to(1+A[[z_{1},z_{2},\ldots,z_{n}]]^{+})^{\times}$
@@
$\operatorname{Exp}\left(\sum_{r_{1},r_{2},\ldots,r_{n}}A_{r_{1},r_{2},\ldots,r_{n}}\prod_{i=1}^{n}z_{i}^{r_{i}}\right)=\prod_{r_{1},r_{2},\ldots,r_{n}}\lambda^{-1}_{A_{r_{1},r_{2},\ldots,r_{n}}}(-\prod_{i=1}^{n}z_{i}^{r_{i}}).$
@@
Step 1. For $n>0$, let $A_{m}$ denote the ideal of $A[[z_{1},z_{2},\ldots,z_{n}]]$ generated by monomials of degree at least $m$, then $A[[z_{1},z_{2},\ldots,z_{n}]]^{+}=A_{1}$ and $\operatorname{Exp}(A_{m})\subset 1+A_{m}$. Thus, $\operatorname{Exp}$ induces a homomorphism
@@
Since $\operatorname{Exp}$ is an isomorphism, we have the inverse isomorphism, which we call plethystic logarithm: $\operatorname{Log}\coloneqq\operatorname{Exp}^{-1}$. In the next proposition, we can verify the basic properties for them.
@@
The first property follows from the definitions of $\operatorname{Exp}$ and of pre-$\lambda$ structures. The second property holds since $\operatorname{Log}$ is the inverse of $\operatorname{Exp}$. ∎
@@ 4.2. Definition of the power structure
In this section, we give the definition of the power structure and explain how to obtain a power structure from a pre-$\lambda$ structure. This is a generalization of [GZLMH13, Definition 1] to the multivariable case.
@@
    $(1+\sum_{i}z_{i})^{m}=1+\sum_{i}mz_{i}+\text{terms of degree at least }2$;
@@
The properties (i), (ii) and (v) follow from the definitions, while properties (iii) and (iv) follow from Proposition 4.2. The property (vii) is true because $\operatorname{Exp}$ commutes with monomial substitutions. Now it remains to prove property (vi). It follows from the definition of $\operatorname{Exp}$ and of pre-$\lambda$ structure that
@@
The power structures corresponding to different pre-$\lambda$ structures might coincide. As in [GZLMH13, Section 3], the power structures defined by motivic zeta function
$\zeta_{M}(t)=1+[S^{1}M]\cdot t+[S^{2}M]\cdot t^{2}+\cdots,$
@@
$\phi_{M}(t)=1+[M]\cdot t+[M_{1}]\cdot t^{2}+\cdots,$
@@ 4.3. Power structures on the rings of motivic classes of schemes and stacks
In Section 4.1 and Section 4.2 we dealt with arbitrary pre-$\lambda$ structures. Now we apply this formalism to the pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$ defined in Section 2.3. Thus, we obtain a power structure on the ring $\overline{\operatorname{Mot}}(k)$. We have a theorem that is a generalization of [FSS18, Lemma 3.8.2] to arbitrary characteristic:
@@
Assume that $M$ is a scheme and $\mathcal{A}_{1},\mathcal{A}_{2},\ldots,\mathcal{A}_{n},\ldots$ are stacks with $\mathcal{A}_{0}=\operatorname{Spec}k$. Put $A(t)=[\mathcal{A}_{0}]+[\mathcal{A}_{1}]t+[\mathcal{A}_{2}]t^{2}+\cdots+[\mathcal{A}_{n}]t^{n}+\ldots$. Then we have
$(A(t))^{[M]}=1+\sum_{k=1}^{\infty}\left\{\sum_{k_{i}\colon\sum ik_{i}=k}\left[\left(\left(M^{\sum_{i}k_{i}}\backslash\Delta\right)\times\prod_{i}\mathcal{A}_{i}^{k_{i}}\right)/\prod_{i}{S_{k_{i}}}\right]\right\}\cdot t^{k},$
where the power structure is given in Proposition 4.4 and $\Delta$ is the “large diagonal” in $M^{\sum_{i}k_{i}}$ which consists of ($\sum_{i}k_{i}$)-tuples of points of $M$ with at least two coinciding ones.
@@
By the proof of [BM15, Lemma 6], we reduce the stack case to the case where $\mathcal{A}_{i}$ are schemes. Writing $B(t)=\operatorname{Log}A(t)$, we just need to check that $\operatorname{Exp}([M]B(t))=\operatorname{Exp}(B(t))^{[M]}$. Now, by [GZLMH04, Remark 2, Page 55],
$\operatorname{Exp}([M]B(t))=\prod_{k\geq 1}(1-t^{k})^{-[M]B(t)},$
@@
Theorem 4.6 shows that the power structure is effective in the following sense: if $M$ is a scheme and the coefficient of $A(t)$ are motivic classes of stacks, then the coefficients of $(A(t))^{[M]}$ are also motivic classes of stacks. This is not always true if $M$ is a stack and there is a counterexample in [GZLMH13, Statement 2]:
$(1+t)^{[\operatorname{BGL}(1)]}=(1+t)^{1/(\mathbb{L}-1)}$
is not effective since the coefficient at $t^{2}$ is equal to

### v2

@@ 3.2. $\lambda$-ring $R_{q}$
takes $F^{m}\operatorname{Mot}_{sch}(\mathbb{F}_{q})[\mathcal{L}^{-1}]$ to $F^{2m}R_{q}$, so it is continuous and can be extended to the completion $\overline{\operatorname{Mot}}_{sch}(\mathbb{F}_{q})$. Finally $\overline{\operatorname{Mot}}_{sch}(\mathbb{F}_{q})\cong\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ by Proposition 2.9. ∎
Recall that $\mathcal{L}\coloneqq[\overline{\mathbb{Q}}_{\ell},q]$ is invertible in $R_{q}$. We define a localization of $R_{q}$ as
$R_{q,fin-type}\coloneqq R_{q}[(\mathcal{L}^{i}-1)^{-1}|i>0].$
Since $\mathcal{L}^{i}-1$ is invertible in $\overline{R}_{q}$, the homomorphism $R_{q}\to\overline{R}_{q}$ factors through $R_{q,fin-type}$. Note that $R_{q}\to\overline{R}_{q}$ is injective, so the induced morphism $R_{q,fin-type}\to\overline{R}_{q}$ is injective as well.
We also define a map $\operatorname{tr}\colon R_{q}\to\overline{\mathbb{Q}}_{\ell}\text{ given by }[V,\psi]\mapsto\operatorname{tr}\psi.$ We claim that it is a ring homomorphism. It suffices to check for the basis $\mathcal{E}_{\lambda}],[\mathcal{E}_{\mu}]\in R_{q}$:
$\operatorname{tr}\left([\mathcal{E}_{\lambda}][\mathcal{E}_{\mu}]\right)=\operatorname{tr}\left([\mathcal{E}_{\lambda\mu}]\right)=\lambda\mu=\operatorname{tr}\left([\mathcal{E}_{\lambda}]\right)\cdot\operatorname{tr}\left([\mathcal{E}_{\mu}]\right).$
Note that the homomorphism $\operatorname{tr}$ takes $\mathcal{L}$ to $q$ and $\mathcal{L}^{i}-1$ to $q^{i}-1\neq 0$, so by the universal property of localizations, it extends to a ring homomorphism $\operatorname{tr}\colon R_{q,fin-type}\to\overline{\mathbb{Q}}_{\ell}.$
@@
We have the following commutative diagram:
[…]
Note that #\# is a homomorphism of Theorem 3.2.
@@
By construction, the image of
@@
is contained in $R_{q,fin-type}$. Thus the upper square of diagram (3.1) commutes.
@@
Recall that $\operatorname{tr}(\mathcal{L})=q$, then by Theorem 2.6 and the universal property of localization, we see that the lower square of diagram (3.1) commutes. ∎
Recall that in Section 3.2 we defined a $\lambda$-ring structure on $R_{q}$. The $\lambda$-ring structure on $R_{q}$ can be extended to $\overline{R}_{q}$ by continuity. We also know that $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ is a pre-$\lambda$-ring by Theorem 2.18.
@@
Let $X$ be a reduced quasi-projective scheme over the finite field $\mathbb{F}_{q}$. Denote by $\lambda_{H}$ the $\lambda$-ring structure on $\overline{R}_{q}$. First we check that
$\lambda_{H}(-H^{\ast}([X]))=H^{\ast}(\lambda(-[X])).$
Set $H^{\ast}(X)\coloneqq\bigoplus_{i}(H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell}),F_{q})$ and view it as a super object of $J_{q}$, so we have $H^{\ast}(X)]\in R_{q}$.
By Lemma 3.5,
$\lambda_{H}(-H^{\ast}([X]))=\sum_{i}[Sym^{i}(H^{\ast}(X))](-t)^{i}.$
On the other hand, we have
$H^{\ast}(\lambda(-[X]))=\sum_{i}H^{\ast}([Sym^{i}X])(-t)^{i}=\sum_{i}[H^{\ast}(Sym^{i}X)](-t)^{i}.$
However by Deligne’s theorem (see [SGA4, Exposé XVII, Théorème 5.5.21]),
$Sym^{i}(H^{\ast}(X))\simeq H^{\ast}(Sym^{i}X).$
Since $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ is topologically generated by classes $\mathbb{L}^{-i}[X]$, the statement follows. ∎
Proof of part (ii) of Theorem 3.2.
@@
Now the statement follows by diagram chasing. ∎
@@ 4. Power structures
Our goal in this section is to define a power structure on $\overline{\operatorname{Mot}}(k)$ and interpret certain powers as configuration spaces, see Proposition 4.4 and Theorem 4.6. These are similar to [GZLMH13, Definition 1] and [BM15, Lemma 6] respectively.
@@ 4.1. From pre-$\lambda$-rings to plethystic exponents and plethystic logarithms
In this section, for an arbitrary pre-$\lambda$-ring, as a generalization of [GZLMH04, Remark 2, Page 55] we define the multivariable generalizations of the plethystic exponents and plethystic logarithms.
Let $R$ be a pre-$\lambda$-ring with the pre-$\lambda$-ring structure $a\mapsto\lambda_{a}(z)$. We denote by $R[[z_{1},z_{2},\ldots,z_{n}]]^{+}$ the ideal of power series with vanishing constant term and denote by $(1+R[[z_{1},z_{2},\ldots,z_{n}]]^{+})^{\times}$ the multiplicative group of series with constant term equal to $1$, we define the plethystic exponent
$\operatorname{Exp}\colon R[[z_{1},z_{2},\ldots,z_{n}]]^{+}\to(1+R[[z_{1},z_{2},\ldots,z_{n}]]^{+})^{\times}$
@@
$\operatorname{Exp}\left(\sum_{r_{1},r_{2},\ldots,r_{n}}A_{r_{1},r_{2},\ldots,r_{n}}\prod_{i=1}^{n}z_{i}^{r_{i}}\right)\coloneqq\prod_{r_{1},r_{2},\ldots,r_{n}}\lambda^{-1}_{A_{r_{1},r_{2},\ldots,r_{n}}}(-\prod_{i=1}^{n}z_{i}^{r_{i}}).$
@@
Step 1. For $n>0$, let $A_{m}$ denote the ideal of $R[[z_{1},z_{2},\ldots,z_{n}]]$ generated by monomials of degree at least $m$, then $R[[z_{1},z_{2},\ldots,z_{n}]]^{+}=A_{1}$ and $\operatorname{Exp}(A_{m})\subset 1+A_{m}$. Thus, $\operatorname{Exp}$ induces a homomorphism
@@
Since $\operatorname{Exp}$ is an isomorphism, we have the inverse isomorphism, which we call plethystic logarithm: $\operatorname{Log}\coloneqq\operatorname{Exp}^{-1}$. In the next proposition, we verify the basic properties for them.
@@
The first property follows from the definitions of $\operatorname{Exp}$ and of pre-$\lambda$-ring structures. The second property holds since $\operatorname{Log}$ is the inverse of $\operatorname{Exp}$. ∎
@@ 4.2. Definition of the power structure
In this section, we give the definition of the power structure and explain how to obtain a power structure from a pre-$\lambda$-ring structure. This is a generalization of [GZLMH13, Definition 1] to the multivariable case.
@@
    $(1+\sum_{i}z_{i})^{m}=1+\sum_{i}mz_{i}+\text{terms of degree at least }2$.
@@
The properties (i), (ii) and (v) follow from the definitions, while properties (iii) and (iv) follow from Proposition 4.2. The property (vii) is true because $\operatorname{Exp}$ commutes with monomial substitutions. Now it remains to prove property (vi). It follows from the definition of $\operatorname{Exp}$ and of pre-$\lambda$-ring structure that
@@
The power structures corresponding to different pre-$\lambda$-ring structures might coincide. As in [GZLMH13, Section 3], the power structures defined by motivic zeta-function
$\zeta_{M}(z)=1+[S^{1}M]\cdot z+[S^{2}M]\cdot z^{2}+\cdots,$
@@
$\phi_{M}(z)=1+[M]\cdot z+[M_{1}]\cdot z^{2}+\cdots,$
@@ 4.3. Power structures on the rings of motivic classes of schemes and stacks
In Section 4.1 and Section 4.2 we dealt with arbitrary pre-$\lambda$-ring structures. Now we apply this formalism to the pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$ defined in Section 2.3. Thus, we obtain a power structure on the ring $\overline{\operatorname{Mot}}(k)$. We have a theorem that is a generalization of [FSS18, Lemma 3.8.2] to arbitrary characteristic.
@@
Assume that $M$ is a scheme, $\mathcal{A}_{0}=\operatorname{Spec}k$ and $\mathcal{A}_{1},\mathcal{A}_{2},\ldots,\mathcal{A}_{n},\ldots$ are stacks. Put $A(z)=[\mathcal{A}_{0}]+[\mathcal{A}_{1}]z+[\mathcal{A}_{2}]z^{2}+\cdots+[\mathcal{A}_{n}]z^{n}+\ldots$. Then we have
$(A(z))^{[M]}=1+\sum_{k=1}^{\infty}\left\{\sum_{k_{i}\colon\sum ik_{i}=k}\left[\left(\left(M^{\sum_{i}k_{i}}\backslash\Delta\right)\times\prod_{i}\mathcal{A}_{i}^{k_{i}}\right)/\prod_{i}{S_{k_{i}}}\right]\right\}\cdot z^{k},$
where the power structure is given in Proposition 4.4 and $\Delta$ is the “large diagonal” in $M^{\sum_{i}k_{i}}$ which consists of $\left(\sum_{i}k_{i}\right)$-tuples of points of $M$ with at least two coinciding ones.
@@
By the proof of [BM15, Lemma 6], we reduce the stack case to the case where $\mathcal{A}_{i}$ are schemes. Writing $B(z)=\operatorname{Log}A(z)$, we just need to check that
$\operatorname{Exp}([M]B(z))=\operatorname{Exp}(B(z))^{[M]}.$
Now, by [GZLMH04, Remark 2, Page 55],
$\operatorname{Exp}([M]B(z))=\prod_{k\geq 1}(1-z^{k})^{-[M]B(z)},$
@@
Theorem 4.6 shows that the power structure is effective in the following sense: if $M$ is a scheme and the coefficient of $A(z)$ are motivic classes of stacks, then the coefficients of $(A(z))^{[M]}$ are also motivic classes of stacks. This is not always true if $M$ is a stack and there is a counterexample in [GZLMH13, Statement 2]:
$(1+z)^{[\operatorname{BGL}(1)]}=(1+z)^{1/(\mathbb{L}-1)}$
is not effective since the coefficient at $z^{2}$ is equal to